Shift summary for Eve 12/13/99

Julie Whitmore (jaws@fnal.gov)
Tue, 14 Dec 1999 00:20:20 -0600 (CST)

Shift Summary for eve shift, 13/dec/99

On Shift:HW,JAWS

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E832 data (special runs not included):0
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time: No beam. Booster feeder problem.

========================================================================
Runs ended this shift:
Run 15214 - LASER_SCAN

Tapes brought to FCC this shift:
(including tapes for online split)
None.

========================================================================
Accesses during shift:
Hogan and Ashkan remove DC4 postamp power from remote control.
Jason and Julie reseat Et cable to channel 1767.
RickK and Herman swap distriminator channel for new 4th accidental counter.

========================================================================
Overview of shift:
No beam.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 HN loads in DCT0s from run 15101.

15:50 JCH and JAWS make access to look at inefficient
CsI dynode channel 1767. Reseating connector
into Et fixes it.

16:00 HN and AA make access and remove DC4 postamp
power from remote control to see if DC4YU RF
noise improves. AA puts DC4OR trigger source
back in previous "transformer location."

16:40 Install CSQ6 constants from run 15211 as
per Ed's instructions.

16:59 "Booster has lost power. FESS investigating."

17:00 RickC and Herman on access to investigate
accidental counter discriminator problem.

17:05 Herman asks that we turn off NM4AN.

17:26 "Booster has lost power. Estimate 8hrs."

17:40 Take long laser scan 15214.

17:45 RickC and Herman out of access. Swap a
discriminator channel for new 4th accidental
counter.

17:50 DN hooks up 2 new bottles of Ar-Eth.

18:00 RickC looks into BOA problem that Leo had
noticed. Rates are ~20% high but are ok.

19:00 Julie sneaks off for a beer. Herman left
to mind the store.

20:00 Julie returns. Too bad beam hasn't.

21:00 Turn on NM4AN. NM2S3 still set for access
mode. NM3S is off.

========================================================================
Follow up Notes on Previous Shift
CsI 2081 has a large gain drift(~5%), keep an eye on it.
Action - install new constants. Need to update reference plot.

HCC channels 2803 and 1767 were both highly inefficient....
Action - reseat connector for 1767. Channel 2803 is ok.

========================================================================
Items to Follow up on Next Shift

Update CsI laser reference plot when beam returns.